### v1

@@ 3 METHODOLOGY
3.1 Effects of level set parameters in 3D printing simulations
In a 3D printing study, ink deposition is characterized by a velocity that changes from the bottom of the bed to the top interface between ink and air. In addition, at the beginning of the simulation, it takes some time for the fluid to be deposited on the bed, and the use of a fixed reinitialization can create some mass conservation errors due to the proximity of the active (moving) and inactive (stationary) interfaces McCaslin and Desjardins (2014). Based on these conditions and the insights provided in the previous section, we analyzed how mass conservation is affected by the level-set parameters. To avoid issues originating from the interaction of the active and inactive interfaces, we computed the cross sections of the strand 2mm2\ $m$ away from the nozzle axis at t=0.5st=0.5\ s.
@@ 3.1.1 Mass conservation analysis
$V_{pl}=A\bar{v}_{x}$ (11)
where $A_{f}$ is the area of the strand. In addition, the inlet flow rate $V_{pl}$ can be directly related to the cross-sectional area of the nozzle $\pi D^{2}/4$ and the speed of the plunger $v_{p}$, which are the actual parameters of the process. We considered $\Delta Z=0.32\ mm$ and $\bar{v}_{x}=20\ mm/s$ to investigate the effects of the level set parameters. In Comminal et al. (2018); Serdeczny et al. (2018a), the authors compared the compactness of the strands obtained in the simulation with idealized cross sections. In contrast, we directly compared the area $A_{f}$ calculated from the mass conservation in Eq. 11 to the following surface area:
@@
for the different parameters of the level set method considered in this study. For $\epsilon$, we evaluated how the change in thickness around the maximum element size $m_{max}$ affects the cross-sectional area of the fluid. In the case of $\gamma$, we proposed to use values around the nozzle printing speed $\bar{v}_{x}$, as this captures the most critical speed change.
3.1.2 Mass conservation analysis for different printing conditions
Once the effects of the level-set parameter have been established, we simulated the printing process for different settings. All simulations considered for this analysis used the best combination of $\epsilon$ and $\gamma$ found in the parametric study. We defined the same changes ($\Delta Z$) and nozzle speeds ($\bar{v}_{x}$) proposed in Serdeczny et al. (2018a) to validate the results with their simulated and experimental data.
3.2 Mesh convergence analysis
We performed a simple mesh refinement study to test the convergence of the solution for the settings $\Delta Z=0.32\ mm$ and $\bar{v}_{x}=20\ mm/s$. The variables used to evaluate the convergence of the simulations were the surface area $A_{s}$ and the maximum pressure on the nozzle $P_{max}$. With these two variables, we account for the effects of the mesh on both the fluid dynamics and the level-set solutions. For the level set solution, we used as a reference the strand area $A_{f}$ explained in the previous subsection. Since the mesh size affects $\epsilon$, we also considered the difference between $A_{s}$ and $A_{f}$:
@@ 3.2 Mesh convergence analysis
A mesh calibrated for fluid dynamics was created, and the element growth rate (1.08), curve factor (0.3), and resolution of narrow regions (0.95) were fixed. Only the maximum and minimum element sizes were refined during the convergence analysis.
@@ 4 RESULTS
All simulations performed in this work were implemented in the HPC cluster of the Eindhoven University of Technology. The jobs run on only one node with one task and 32 CPUs per task.
4.1 Mesh convergence results
The mesh convergence results are presented in Table II. As a starting point for this analysis, we considered the recommended $\epsilon$ in COMSOL:
where $m_{max}$ and $m_{min}$ are the maximum and minimum element size around the fluid’s interface. We used $\gamma=\{0.02,1\}\ m/s$, the first value is based on the nozzle speed defined for this study, and the second is the default value for the reinitialization parameter in COMSOL.
@@ 4.1 Mesh convergence results
The maximum pressure in the domain $P_{max}$ converged faster than $A_{s}$, which required a finer mesh to reach a relatively convergent value. Although $A_{s}$ was still changing, we did not refine it further because the value obtained for mesh 3 is almost the same as the analytical value $A_{f}$ with less than $1\%$ error. Similarly, $P_{max}$ was less than $5\%$ higher than the reference $\Delta P$. The effects of $\gamma$ are discussed in the following subsection; however, we already notice a big difference between $A_{s}$ and the reference area $A_{f}=62.83e-3\ [mm^{2}]$ (for $\bar{v}_{x}=20\ mm/s$) depending on the value of $\gamma$. Based on the computational time required for $\gamma=0.02$, we decided to use mesh 1 because it took less simulation time compared to mesh 3 and still gave an acceptable error of around $5\%$ in $A_{s}$.
4.2 Effects of the reinitialization parameter
We simulated the process for different values of the reinitialization parameter ($\gamma$) to evaluate its effects on mass conservation. In Fig. 2, we see that the strand cross-section has an ellipsoidal shape and when $\gamma$ increases, the major axis shrinks, while the minor axis expands. The surface areas ($A_{s}$) obtained in the simulation for $\gamma$ values around the nozzle speed $\gamma_{ref}=0.02$ are shown in Table III. Decreasing $\gamma$ to $0.005$ produced a slight improvement from around $5$ to $4\%$ error. In contrast, increasing $\gamma$ had a greater impact on the precision of the results, as can be seen in the errors ($\Delta A$) of the last two parameters in Table III. The simulation time was almost the same for the first two parameters compared to $\gamma_{ref}$. Furthermore, it was also noticed that for the last two parameters considered, there is a significant increase in the computation time.
@@ 4.2 Effects of the reinitialization parameter
4.3 Effects of the interface thickness
Since the default interface thickness used in COMSOL considers the maximum element size, we analyzed the effects of reducing its value. The default interface thickness for mesh 1 was $\epsilon_{ref}=0.043\ [mm]$, which is approximately 7 times less than the nozzle-to-bed distance defined ($\Delta Z=0.32\ [mm]$). For this analysis, we considered changes based on the factor $\epsilon_{f}=\epsilon/\epsilon_{ref}$. Decreasing $\epsilon_{f}$ reduced the mass conservation error, as shown in Table IV. The parameters $\epsilon_{f}=\{0.9,0.8\}$ approximately halve the error, and the following two parameters $\epsilon_{f}=\{0.7,0.6\}$ give an error below $1\%$. Another interesting aspect of the results was that $A_{s}$ starts to exceed the reference value $A_{f}$ for the last two parameters, and there was an increase in the error when reducing $\epsilon_{f}$ from $0.6$ to $0.5$. The qualitative effects of $\epsilon$ on the cross-section are given in Fig. 3. Reducing the factor $\epsilon_{f}$ lead to ellipsoids with a larger major axis; however, the changes in the minor axis did not seem significant. Regarding the computational cost, most parameters resulted in a similar simulation time, except for the last, which took slightly longer than the others.
@@ 4.3 Effects of the interface thickness
Based on the effects of the interface thickness and the influence of the mesh size on $A_{s}$, we also performed an additional test in which we decreased the mesh refinement and found a $\epsilon_{f}$ resulting in a small mass conservation error. The results of using a coarser mesh with $\epsilon_{ref}=0.069\ [mm]$ for different factors $\epsilon_{f}$ are given in Table V. We observed a significant decrease in the cross-sectional error $\Delta A$ from $14$ to $0.2\%$.
@@
The interface thickness for the best result obtained in Table IV is $\epsilon_{4}=0.0258[mm]$, and the best result shown in Table V corresponds to $\epsilon_{5}=0.0276[mm]$. So, what seems to be important is to have an adequate interface thickness around the fluid interface. In terms of the qualitative effects on the cross-section, we compared the two results in Fig. 4. We see that there are some slight differences in the ellipsoids, especially towards the top-centered part of the material. Finally, in terms of computational time, the simulation with $\epsilon_{4}$ took around 6h while the one with $\epsilon_{5}$ took around 3h.
Figure 4: Qualitative differences between $\epsilon_{4}$ and $\epsilon_{5}$
@@ 4.4 Mass conservation analysis for different conditions
To conclude with the numerical analysis, we used the best level set parameters found in the previous subsections ($\gamma=0.005\,\text{m/s},\ \epsilon=0.0276\,\text{mm}$) to evaluate the precision of the simulations for different nozzle-to-bed distances ($\Delta Z$) and nozzle speeds ($\bar{v}_{x}$). The reference areas were scaled according to the ratio $\bar{v}_{x}/v_{p}$ and the results are summarized in Table VI.
@@
Figure 5: Cross-sectional shapes - $\Delta Z/D=1.0$
Figure 6: Cross-sectional shapes - $\Delta Z/D=0.8$
We see that independently of the printing parameters, the errors are very small. In Figures 5 to 7, we present the cross-sectional areas obtained for these simulations. Although the reference areas $A_{f}$ did not change with $\Delta Z/D$, the major axis of the strand shrank when the nozzle-to-bed distance was increased. We also observed that the qualitative behavior of the results was good compared to the experimental results given in Serdeczny et al. (2018a).
Table 7: Aspect ratio comparison with reference
@@
In Table VII, we present the aspect ratio $W/H$ obtained in this work and the values reported in Serdeczny et al. (2018a). There is a good agreement between the simulated (Sim) and experimental (Exp) data in Serdeczny et al. (2018a) and the results obtained in this work for almost all the conditions evaluated. Furthermore, the shapes of the areas seem to match the experimental results more accurately than the areas simulated in Serdeczny et al. (2018a).
Figure 7: Cross-sectional shapes - $\Delta Z/D=0.6$

### v2

@@ 3 METHODOLOGY
3.1 Simulation settings
The simulations were implemented in COMSOL Multiphysics® v6.2 (comsol), using the Laminar Flow and level-set interfaces coupled with the Two-Phase Flow level-set multiphysics. A stationary study step was defined to have a consistent initialization of the time-dependent solution, and the boundary conditions were ramped up to avoid numerical issues. All simulations performed in this work were implemented in the HPC cluster of the Eindhoven University of Technology. The jobs run on only one node with one task and 32 CPUs per task.
3.2 Parametric analysis of the level-set method
In the present study, we assess the influence of the level-set parameters on the numerical results. As presented in Section 2.2, the level-set parameters affect the numerical accuracy and the computational cost of the CLS solutions. In addition, the analytical relationships available for $\epsilon$ and $\gamma$ are only valid for specific discretization schemes. Therefore, we propose a parametric analysis to isolate the individual effects of $\epsilon$ and $\gamma$ on the mass conservation of 3D printing simulations. For simplicity, this analysis is done considering single-layer simulations, but is evaluated with multilayer printing scenarios.
@@ 3.2.1 Mass conservation analysis
$V_{pl}=A_{f}\bar{v}_{x}$ (13)
where $A_{f}$ is the area of the strand and $\bar{v}_{x}$ is the printing speed. In addition, the inlet flow rate $V_{pl}$ can be directly related to the cross-sectional area of the nozzle $\pi D^{2}/4$ and the speed of the plunger $v_{p}$, which are settings of the process. We considered $\Delta Z=0.32\ mm$ and $\bar{v}_{x}=20\ mm/s$ to investigate the effects of the level-set parameters. In 1_comminal_numerical_2018; 2_serdeczny_experimental_2018, the authors compared the compactness of the strands obtained in the simulation with idealized cross sections. In contrast, we directly compared the area $A_{f}$ calculated from the mass conservation in Eq. 13 to the following surface area:
@@
with $\xi$ evaluated on all elements and $A_{s}$ calculated on a predefined cut plane for the different parameters of the level-set method considered in this study. For $\epsilon$, we evaluated how the change in thickness around the maximum mesh element size $m_{max}$ affects the cross-sectional area of the fluid. In the case of $\gamma$, we proposed to use values around the printing speed $\bar{v}_{x}$, as this captures the most critical speed change. At the beginning of the simulation, the fluid takes some time to spread over the bed, and a fixed reinitialization rate can lead to mass conservation errors due to the proximity of the active (moving) and inactive (stationary) interfaces (14_mccaslin_localized_2014). Therefore, the mass conservation analysis was carried out using cross-sections of the strand 2mm2\ $\mathrm${mm} away from the nozzle axis at $t=0.5\mathrm{s}$.
3.2.2 Mesh convergence analysis
We performed a simple mesh refinement study to test the convergence of the solution for the settings $\Delta Z=0.32\ mm$ and $\bar{v}_{x}=20\ mm/s$. As a starting point for this analysis, we considered the recommended $\epsilon$ in COMSOL:
where $m_{max}$ and $m_{min}$ are the maximum and minimum element size around the fluid’s interface. We used $\gamma=\{0.02,1\}\ m/s$, the first value is based on the printing speed defined for this study, and the second is the default value for the reinitialization parameter in COMSOL. The variables used to evaluate the convergence of the simulations were the surface area $A_{s}$ and the maximum pressure on the nozzle $P_{max}$. With these two variables, we account for the effects of the mesh on both the fluid dynamics and the level-set solutions. For the level-set solution, we used as a reference the strand area $A_{f}$ given by Eq. 13. Since the mesh size affects $\epsilon$, we also considered the difference between $A_{s}$ and $A_{f}$:
@@ 3.2.2 Mesh convergence analysis
A physics-controlled mesh with a fine element size calibrated for fluid dynamics was defined. In the mesh settings, only the maximum and minimum element sizes were changed during the convergence analysis. The other mesh settings, element growth rate (1.08), curve factor (0.3), and resolution of narrow regions (0.95), were kept constant in all the numerical solutions.
@@ 4 RESULTS
In this section, we present a parametric study to evaluate the accuracy of the conservative level-set formulation as a function of the reinitialization factor and interface thickness. The analysis focuses on assessing mass conservation and examining the resulting strand geometry over a range of level-set parameters selected based on recommendations from the literature. Building upon this framework, we further extend the simulations to multilayer printing scenarios involving Bingham fluids with different rheological properties and nozzle cross-section geometries.
4.1 Effects of the level-set parameters in single-layer printing
This subsection investigates the sensitivity of the simulated strand geometry and mass conservation to variations in the interface thickness and the reinitialization parameters within the conservative level-set framework. We begin by assessing mesh convergence using the ideal strand area derived from the mass balance and the maximum pressure within the domain. This is followed by a systematic set of single-layer deposition tests aimed at quantifying the influence of the level-set parameters on the accuracy of the numerical results.
4.1.1 Mesh convergence results
The values of reference for the mesh convergence test were $A_{f}=62.83\times 10^{-3}\ [mm^{2}]$ (for $\bar{v}_{x}=20\ mm/s$) and $\Delta P=8$ [MPa]. The convergence results are presented in Table 3.
@@ 4.1.1 Mesh convergence results
The maximum pressure in the domain $P_{max}$ converged faster than $A_{s}$, which required a finer mesh to reach a relatively convergent value. Although $A_{s}$ was still changing, we did not refine it further because the value obtained for Mesh 3 is almost the same as the analytical value $A_{f}$ with less than $1\%$ error. Similarly, $P_{max}$ exceeded the reference $\Delta P$ by less than $5\%$. The effects of $\gamma$ are discussed in the following subsection; however, we already notice a big difference between $A_{s}$ and the reference area $A_{f}$ depending on the value of $\gamma$. Based on the computational time required for $\gamma=0.02$, we decided to use Mesh 1 because it took less simulation time compared to Mesh 3 and still gave an acceptable error of around $5\%$ in $A_{s}$.
4.1.2 Effects of the reinitialization parameter
We simulated the process for different values of the reinitialization parameter ($\gamma$) to evaluate its effects on mass conservation. In Fig. 3, we see that the strand cross-section has an ellipsoidal shape and when $\gamma$ increases, the major axis shrinks, while the minor axis expands. The surface areas ($A_{s}$) obtained in the simulation for $\gamma$ values around the printing speed $\gamma_{ref}=0.02\ m/s$ are shown in Table 4. Decreasing $\gamma$ to 0.005m/s0.005\ m/s produced a slight improvement from around $5$ to $4\%$ error. In contrast, increasing $\gamma$ had a greater impact on the precision of the results, as can be seen in the errors ($\Delta A$) of the last two parameters $\gamma=\{0.04,0.08\}\ m/s$ in Table 4. The simulation time was almost the same for the first two parameters compared to $\gamma_{ref}$. Furthermore, it was also noticed that for the last two parameters considered, there is a significant increase in the computation time.
@@ 4.1.2 Effects of the reinitialization parameter
4.1.3 Effects of the interface thickness
Since the default interface thickness used in COMSOL considers the maximum element size, we analyzed the effects of reducing its value. The default interface thickness for Mesh 1 was $\epsilon_{ref}=0.043\ [mm]$, which is approximately 7 times less than the nozzle-to-bed distance defined ($\Delta Z=0.32\ [mm]$). For this analysis, we considered changes based on the factor $\epsilon_{f}=\epsilon/\epsilon_{ref}$ and used $\gamma=0.02\ m/s$. Decreasing $\epsilon_{f}$ reduced the mass conservation error, as shown in Table 5. The parameters $\epsilon_{f}=\{0.9,0.8\}$ approximately halve the error, and the following two parameters $\epsilon_{f}=\{0.7,0.6\}$ give an error below $1\%$. Another interesting aspect of the results was that $A_{s}$ starts to exceed the reference value $A_{f}$ for the last two parameters, and there was an increase in the error when reducing $\epsilon_{f}$ from $0.6$ to $0.5$. The qualitative effects of $\epsilon$ on the cross-section are given in Fig. 4. Reducing the factor $\epsilon_{f}$ lead to ellipsoids with a larger major axis; however, the changes in the minor axis did not seem significant. Regarding the computational cost, most parameters resulted in a similar simulation time, except for the last, which took slightly longer than the others.
@@ 4.1.3 Effects of the interface thickness
Based on the effects of the interface thickness and the influence of the mesh size on $A_{s}$, we also performed an additional test in which we decreased the mesh refinement and found a $\epsilon_{f}$ resulting in a low mass conservation error. The results of using a coarser mesh with $\epsilon_{ref}=0.069\ [mm]$ for different factors $\epsilon_{f}$ are given in Table 6. We observed a significant decrease in the cross-sectional error $\Delta A$ from $14$ to $0.2\%$.
@@
The interface thickness for the best result obtained in Table 5 ($\epsilon_{f}=0.6$) is $\epsilon=0.0258[mm]$, and the best result shown in Table 6 ($\epsilon_{f}=0.4$) corresponds to $\epsilon=0.0276[mm]$. Therefore, what seems to be important is to have an adequate interface thickness around the fluid interface. In terms of the qualitative effects on the cross-section, we compared the two results in Fig. 5. We see that there are some slight differences in the ellipsoids, especially towards the top-centered part of the material. Finally, in terms of computational time, the simulation with the finer mesh and $\epsilon_{f}=0.6$ took around 6h while the one with the coarser mesh and $\epsilon_{f}=0.4$ took around 3h.
Figure 5: Cross-sections for different mesh refinements with similar interface thickness
@@ 4.2 Mass conservation analysis for different conditions
To conclude with the numerical analysis of single layer printing, we used the best level-set parameters found in the previous subsections ($\gamma=0.005\,\text{m/s},\ \epsilon=0.0276\,\text{mm}$) to evaluate the precision of the simulations for different nozzle-to-bed distances ($\Delta Z$) and printing speeds ($\bar{v}_{x}$). The reference areas were scaled according to the ratio $\bar{v}_{x}/v_{p}$ and the results are summarized in Table 7.
@@
We see that independently of the printing parameters, the errors ($\Delta A$) are very low. In Figure 6, we present the cross-sectional areas obtained for these simulations. Although the reference areas $A_{f}$ did not change with $\Delta Z/D$, the major axis of the strand shrank when the nozzle-to-bed distance was increased. We also observed that the qualitative behavior of the results was good compared to the experimental results given in 2_serdeczny_experimental_2018.
Table 8: Aspect ratio comparison with W/HSimW/H-Sim and W/HExpW/H-Exp in 2_serdeczny_experimental_2018
@@
In Table 8, we present the aspect ratio $W/H$ obtained in this work and the values reported in 2_serdeczny_experimental_2018. There is a good agreement between the simulated (Sim) and experimental (Exp) data in 2_serdeczny_experimental_2018 and the results obtained in this work for almost all the conditions evaluated. Furthermore, the shapes of the areas seem to match the experimental results more accurately than the areas simulated in 2_serdeczny_experimental_2018.
Figure 6: Cross-sectional shapes for different printing conditions
4.3 Two-layer printing of viscoplastic fluids
In this subsection, we extend the previous analysis to evaluate the influence of the level-set parameters on multilayer printing involving varying fluid rheologies and nozzle geometries. In general, multilayer printing involves a complex interplay between layers, as a previously printed layer deforms when a new layer is deposited on top. Different printing strategies are commonly employed to achieve stable printing without excessive deformation. These include (i) wet-on-wet (11_mollah_stability_2021) (ii) wet-on-semisolid (431_mollah_computational_2023), (iii) wet-on-solidifying (431_mollah_virtual_2025), and (iv) wet-on-solid (11_mollah_stability_2021) approaches, particularly when printing viscoplastic materials such as thermosets, cementitious materials, and geopolymers. In this study, we focus on the wet-on-wet printing strategy, where the material rheology does not change during the printing process. Viscoplastic materials with different rheological properties are considered, together with nozzles featuring cylindrical and rectangular cross-sections. The analysis focuses on assessing how the level-set parameters affect strand morphology and overall mass conservation, and the results obtained with the conservative level-set formulation are compared against validated data from the literature.
4.3.1 Concrete printing
We selected the level-set parameters, $\epsilon$ and $\gamma$, following the insights gained from the results described in section 4.1. The reinitialization was defined as a fraction of the horizontal speed, $\gamma=0.25|\bar{v}_{x}|$, and the interface thickness as a function of the maximum element size $\epsilon=0.3m_{max}$. The maximum and minimum element sizes for the mesh were defined in terms of the diameter of the nozzle, $m_{min}=0.01D$ and $m_{max}=0.1D$. The other mesh settings were defined by the physics-based mesh feature in COMSOL.
To assess the accuracy of the simulation, we examined the strand cross sections for the first and second layers, as shown in Figure 7. The cross section of the first layer was obtained by defining a cut plane at $x=200\,\text{mm}$ at $t=9\,\text{s}$. For the second layer, we tracked the motion of this reference section from $t=9\,\text{s}$ to $t=20\,\text{s}$, which, based on the imposed printing speed, corresponded to a position of $x=424.4\,\text{mm}$. According to the values in Table 2, the reference cross-sectional area for a single layer is $A_{f}=550\,\text{mm}^{2}$, giving $A_{f}=1100\,\text{mm}^{2}$ for two layers.
In Figure 7, we compare the results obtained in this work (solid line) with the experimental measurements reported by 10_spangenberg_numerical_2021, as well as the numerical results obtained using the volume of fluid method (VOF) combined with a generalized Newtonian model, as presented in el_abbaoui_3d_2024. At $t=9\,\text{s}$, the contour plot shows a strand width of approximately $46\,\text{mm}$, which is approximately $5\,\text{mm}$ wider than the measurements and slightly larger than the VOF results with both no-slip (NS) and free slip (FS) boundary conditions at the nozzle wall. The simulated cross-sectional area was $A_{s}=548\,\text{mm}^{2}$, corresponding to an error ($\Delta A$) of $0.36\%$. For the second layer, we observed a slight increase in the width of the first-layer strand to about $48\,\text{mm}$, along with minor changes in its contour. The upper strand exhibited a width of roughly $44\,\text{mm}$ and appeared less rounded than the bottom strand. The computed area for the two-layer configuration was $A_{s}=1097\,\text{mm}^{2}$, yielding an error ($\Delta A$) of $0.27\%$.
Overall, the two-layer cross-section obtained in this work showed good agreement with both the experimental data and the reference simulations. However, the interface between the two layers is smoothed when using the level-set method. As explained in the model description, with the level-set method, the material boundaries are defined by a diffuse and constant interface of finite thickness, and the properties of the phases are regularized with a Heaviside function that transitions smoothly between phases across multiple mesh elements. Consequently, when multiple layers are deposited sequentially, previously formed interfaces are continuously advected and reinitialized, resulting in numerical blending of layers rather than preservation of distinct material boundaries.
Among the reference results, the VOF simulations with free-slip (FS) boundary conditions were closest to the experimental measurements. Despite the use of a simplified constitutive model and possible variations in the vertical motion profile between layers, the present results remain consistent with the reference studies (10_spangenberg_numerical_2021; el_abbaoui_3d_2024) and with the analysis of mass-conservation.
Figure 7: Cross-sectional shapes for multilayer concrete simulations
4.3.2 Metakaolin printing
A similar analysis to the one described in the previous subsection was performed for the Metakaolin printing, in which we again evaluated the strand cross-sections for the first and second layers. The resulting cross-sections are shown in Figure 8. The profile of the first layer was obtained using a cutting plane at x=20mmx=-20\ $\mathrm${mm} at t=9st=9\ $\mathrm${s}. For the second layer, we tracked the expected motion of this cutting plane according to the prescribed printing speed, reaching x=20mmx=20\ $\mathrm${mm} at t=22st=22\ $\mathrm${s}. Based on the reference values listed in Table 2, the theoretical cross-sectional area for a single layer is $A_{f}=36\ \mathrm{mm}^{2}$, hence $A_{f}=72\ \mathrm{mm}^{2}$ for two layers. Finally, the simulated cross-sections were also compared with the experimental results reported in Mollah2025RapidCuringCFD.
Figure 8: Cross-sectional shapes for multilayer Metakaolin simulations
The contour plot for the first layer gives a strand width of w=21.2mmw=21.2\ $\mathrm${mm}, which is in good agreement with both the numerical results and the experimental measurements reported in Mollah2025RapidCuringCFD, where the width was approximately w=23mmw=23\ $\mathrm${mm}. The simulated cross-sectional area for the first layer was $A_{s}=36.1\ \mathrm{mm}^{2}$, corresponding to an error of only 0.28% relative to the reference area $A_{f}$. Slight differences were observed in the width (1 mm), and the height (0.3 mm) when comparing the shape of the simulated (solid line) and measured cross-sections, shown in Figure 8. The simulated cross-sectional area of the two layers was $A_{s}=73.85\ \mathrm{mm}^{2}$, corresponding to a deviation of 2.6% from the reference area $A_{f}$. Since the areas based on mass conservation are accurate, the discrepancies in first-layer width may be due to surface tension effects, the model regularization (Bingham-Papanastasiou), material slip on the substrate, or limitations in measurement resolution. Regarding the two-layer profile, the width of the material in the first layer increased slightly to approximately 22.5mm22.5\ $\mathrm${mm}, while the strand of the second layer was roughly 22mm22\ $\mathrm${mm}. The width of the first layer was therefore slightly lower than the experimental value of w=24mmw=24\ $\mathrm${mm} reported in Mollah2025RapidCuringCFD, while the width of the second layer closely matched the measurements reported. Furthermore, we also observed a smooth contrast between the two deposited layers, which is a consequence of the finite thickness interface representation of the conservative level-set method, as explained in the previous subsection.